OpenStudy (volleyballlover55):

Simplify eight to the negative fifth divided by eight to the negative third. one divided by the quantity eight squared one divided by the quantity eight to the eighth power 8^2 8^8

OpenStudy (volleyballlover55):

\[\frac{ 8^{-5} }{ 8^{-3} }\]

OpenStudy (volleyballlover55):

@FaiqRaees

OpenStudy (faiqraees):

\[\large\rm \frac{a^{x}}{a^{y}}= a^{x-y} \]

OpenStudy (volleyballlover55):

8^-8 @FaiqRaees

OpenStudy (volleyballlover55):

which is also considered to be b

OpenStudy (faiqraees):

No (Hint: The power in the denominator is NEGATIVE)

OpenStudy (volleyballlover55):

theres no option that the power is negative

OpenStudy (faiqraees):

\[\large\rm \frac{8^{-5}}{8^{-3}}= 8^{-5-(-3)} \]

OpenStudy (volleyballlover55):

-2? @FaiqRaees

OpenStudy (faiqraees):

8^-2 = 1/8^2

OpenStudy (volleyballlover55):

ok